# The computational power of discrete chemical reaction networks with bounded executions

### v2

@@ 2.4 Time model
The following model of stochastic chemical kinetics is widely used in quantitative biology and other fields dealing with chemical reactions between species present in small counts [Gillespie77]. It ascribes probabilities to execution sequences, and also defines the time of reactions, allowing us to study the computational complexity of the CRN computation in Sections 4 and 5. If the volume is defined to be $n$, the total number of molecules, then the time model is essentially equivalent to the notion of parallel time studied in population protocols [AngluinAE2008Fast]. In this paper, the rate constants of all reactions are $1$, and we define the kinetic model with this assumption. A reaction is unimolecular if it has one reactant and bimolecular if it has two reactants. We use no higher-order reactions in this paper.
@@
The kinetic model is based on the physical assumption of well-mixedness that is valid in a dilute solution. Thus, we assume the finite density constraint, which stipulates that a volume required to execute a CRN must be proportional to the maximum molecular count obtained during execution [SolCooWinBru08]. In other words, the total concentration (molecular count per volume) is bounded. This realistically constrains the speed of the computation achievable by CRNs.
@@ 2.5 Semilinear sets, predicates, functions
The following is a famous characterization of the computational power of CRNs [Angluin2006ComputationalPower, chen2023rate].
Theorem 2.3 ([Angluin2006ComputationalPower, chen2023rate]).
@@ 3 Execution bounded chemical reaction networks
This is a distinct concept from the notion of “bounded” CRNs studied by Rackoff [Rackoff1978CoveringBoundedness] (studied under the equivlaent formalism of vector addition systems). That paper defines a CRN to be bounded from a configuration $\mathbf{x}$ if $|\mathsf{reach}(\mathbf{x})|$ is finite (and shows that the decision problem of determining whether this is true is $\mathsf{EXPSPACE}$-complete.) We use the term execution bounded to avoid confusion with this concept.
@@
We first observe that being execution bounded from $\mathbf{x}$ implies a slightly stronger condition: there is a uniform upper bound on the length of all executions from $\mathbf{x}$111 In other words, this rules out the possibility that, although all executions from $\mathbf{x}$ are finite, there are infinitely many of them $\mathcal{E}_{1},\mathcal{E}_{2},\dots$, each longer than the previous.
@@
A CRN is execution bounded from $\mathbf{x}$ if and only if there is a constant $N\in\mathbb{N}$ such that all executions from $\mathbf{x}$ have length at most $N$. Equivalently, there are finitely many executions from $\mathbf{x}$.
@@
A CRN is execution bounded from $\mathbf{x}$ if and only if $\mathsf{reach}(x)=\{\mathbf{y}\mid\mathbf{x}\Rightarrow\mathbf{y}\}$ is finite and, for all $\mathbf{y}\in\mathsf{reach}(x)$, $\mathbf{y}\not\Rightarrow\mathbf{y}$ except by the zero-length execution.
@@
Every configuration reachable from $\mathbf{x}$ is reached through some execution contained in $T_{\mathcal{C}}(\mathbf{x})$ as a node and there exists only a finite number of them (Section 3). Multiple unique executions can produce the same configuration but one execution cannot produce multiple configurations. Thus, there exists a surjection from the nodes of $T_{\mathcal{C}}(\mathbf{x})$ into $\mathsf{reach}(x)$ and $\mathsf{reach}(x)$ must also be finite. For the second part of the condition, we prove its contrapositive and assume there exists $\mathbf{y}\in\mathsf{reach}(x)$, $\mathbf{y}\Rightarrow_{P}\mathbf{y}\land|P|>0$. Let $P=(\mathbf{p}_{1},\mathbf{p}_{2},\dots,\mathbf{p}_{n})$. It holds that $\mathbf{p}_{1}=\mathbf{p}_{n}$ and $\mathbf{p}_{n-1}\Rightarrow\mathbf{p}_{1}$. We can construct an infinite-length execution $P^{\prime}=(\mathbf{x},\dots,\mathbf{p}_{1},\dots,\mathbf{p}_{n-1},\mathbf{p}_{1},\dots)$, which must also be a valid under the reactions of $\mathcal{C}$, making $\mathcal{C}$ execution unbounded from $\mathbf{x}$.
If $\mathsf{reach}(x)$ is finite and contains no such $\mathbf{y}$, then we can construct a finite, directed, acyclic graph $G_{\mathcal{C}}(\mathbf{x})=(V,E)$ where $V=\mathsf{reach}(x)$, $E=\{(\mathbf{x},\mathbf{y})\mid\mathbf{x}\Rightarrow_{P}\mathbf{y}\land|P|>0\}$. The longest path in the graph has length of at most $|\mathsf{reach}(x)|-1$. A bijection exists between paths in $G_{\mathcal{C}}(\mathbf{x})$ and executions possible in $\mathcal{C}$ starting from $\mathbf{x}$. We set $n=|\mathsf{reach}(x)|$ satisfying that each execution has length of at most $n$, making $\mathcal{C}$ execution bounded.
@@
A execution $\mathcal{E}=(\mathbf{x}_{1},\mathbf{x}_{2},\dots)$ is self-covering if for some $i<j$, $\mathbf{x}_{i}\leqq\mathbf{x}_{j}$. It is strictly self-covering if $\mathbf{x}_{i}\leq\mathbf{x}_{j}$. We also refer to these as (strict) self-covering paths.222 Rackoff [Rackoff1978CoveringBoundedness] uses the term “self-covering” to mean what we call strictly self-covering here, and points out that Karp and Miller [karp1969parallel] showed that $|\mathsf{reach}(\mathbf{x})|$ is infinite if and only if there is a strictly self-covering path from $\mathbf{x}$. The distinction between these concepts is illustrated by the CRN $A\mathop{\rightleftharpoons}\limits B$. From any configuration $\mathbf{x}$, $\mathsf{reach}(\mathbf{x})$ is finite ($|\mathsf{reach}(\mathbf{x})|=\mathbf{x}(A)+\mathbf{x}(B)+1$), and there is no strict self-covering path. However, from (say) $\{A\}$, there is a (nonstrict) self-covering path $\{A\}\Rightarrow\{B\}\Rightarrow\{A\}$, and by repeating, this CRN has an infinite cycling execution within its finite configuration space $\mathsf{reach}(\{A\})=\{\{A\},\{B\}\}$.
@@
For the forward direction, assume there is a self-covering path from $\mathbf{x}$, which reaches to $\mathbf{x}_{i}$ and later to $\mathbf{x}_{j}\geqq\mathbf{x}_{i}$. Then the reactions going from $\mathbf{x}_{i}$ to $\mathbf{x}_{j}$ can be repeated indefinitely (in a cycle if $\mathbf{x}_{i}=\mathbf{x}_{j}$, and increasing some molecular counts unboundedly if $\mathbf{x}_{i}\leq\mathbf{x}_{j}$), so $\mathcal{C}$ is not execution bounded from $\mathbf{x}$.
@@ 4 Execution bounded CRDs stably decide all semilinear sets
In this section, we will show the computational equivalence between execution bounded and execution unbounded CRNs by a construction. The following is the main result of this section.
@@
Every mod set $M=\big{\{}\left(x_{1},\ldots,x_{d}\right)\mid\sum_{i=1}^{d}w_{i}x_{i}\equiv c\bmod m\big{\}}$ is stably decidable by an execution bounded, single-voting CRD with stabilization time $\Theta(n\log n)$.
@@
Every threshold set $T=\big{\{}(x_{1},\ldots,x_{d})\mid\sum_{i=1}^{d}w_{i}x_{i}\geq t\big{\}}$ is stably decidable by an execution bounded, single-voting CRD with stabilization time $\Theta(n\log n)$.
@@
If sets $X_{1},X_{2}\subseteq\mathbb{N}^{d}$ are stably decided by some execution bounded, single-voting CRD, then so are $X_{1}\cup X_{2},X_{1}\cap X_{2}$, and $\overline{X_{1}}$ with stabilization time $O(n\log n)$.
@@
by two separate CRDs. Subsequently, we add reactions to record the separate votes in one of four new voter species: $V_{NN},V_{NY},V_{YN},V_{YY}$. The first and second CRN determine the first and second subscript respectively. For $b\in\{Y,N\}$ and if $S_{b},T_{b}$ are leaders of $\mathcal{C}_{1}$ and $\mathcal{C}_{2}$ respectively, add the reactions:
@@
E.g. if $N_{1}$ is the no voter of the first CRD, we would add $N_{1}+L_{YN}\rightarrow N_{1}+L_{NN}$ and $N_{1}+L_{YY}\rightarrow N_{1}+L_{NY}$. We let the yes voters be: $\Upsilon=\{V_{NY},V_{YN},V_{YY}\}$ to stably decide $X_{1}\cup X_{2}$ or $\Upsilon=\{V_{YY}\}$ to stably decide $X_{1}\cap X_{2}$.
@@
Since semilinear predicates are exactly Boolean combinations of threshold and mod predicates, Sections 4, 4 and 4.5 imply the following.
Theorem 4.7.
Every semilinear set is stably decidable by an execution bounded, single-voting CRD, with stabilization time $O(n\log n).$
@@
Every semilinear set is stably decidable by an execution bounded, all-voting CRD, with stabilization time $O(n\log n).$
Proof 4.8.
@@ 5 Execution bounded CRCs stably compute all semilinear functions
Similar to [Chen2012DeterministicFunction, Doty2013Leaderless], we compute semilinear functions by decomposing them into “affine pieces”, which we will show can be computed by execution bounded CRNs and combined by using semilinear predicates to decide which linear function to apply for a given input.333 While this proof generalizes to multivariate output functions as in [Chen2012DeterministicFunction, Doty2013Leaderless], to simplify notation we focus on single output functions. Multi-valued functions $f:\mathbb{N}^{d}\to\mathbb{N}^{l}$ can be equivalently thought of as $l$ separate single output functions $f_{i}:\mathbb{N}^{d}\to\mathbb{N}$, which can be computed in parallel by independent CRCs.
@@
We say that a partial function $f:\mathbb{N}^{k}\rightarrow\mathbb{N}^{2}$ is a diff-representation of $f$ if $\operatorname{dom}f=\operatorname{dom}\hat{f}$ and, for all $\mathbf{x}\in\operatorname{dom}f$, if $\left(y_{P},y_{C}\right)=\hat{f}(\mathbf{x})$, then $f(\mathbf{x})=y_{P}-y_{C}$, and $y_{P}=O(f(\mathbf{x}))$. In other words, $\hat{f}$ represents $f$ as the difference of its two outputs $y_{P}$ and $y_{C}$, with the larger output $y_{P}$ possibly being larger than the original function’s output, but at most a multiplicative constant larger [Doty2013Leaderless].
@@
Let $f:\mathbb{N}^{k}\rightarrow\mathbb{N}$ be an affine partial function. Then there is a diff-representation $\hat{f}:\mathbb{N}^{k}\longrightarrow\mathbb{N}^{2}$ of $f$ and an execution bounded CRC that monotonically stably computes $\hat{f}$ in expected time $O(n)$.
@@
For the $c_{i}$ offset, we must reduce the number of $X_{i}$ by $c_{i}$. Since the result will be used in the next reaction, we want to produce a new species $X_{i}^{\prime}$ and require $X_{i}^{\prime}$ to not be consumed during the computation. We achieve this by adding reactions that let $X_{i}$ consume itself $c_{i}$ times (keeping track with a subscript) and converting $X_{i}$ to $X_{i}^{\prime}$ once $c_{i}$ has been reached. For each $i\in\{1,\ldots,k\}$ and $m,p\in\left\{1,\ldots,c_{i}\right\}$, if $m+p\leq c_{i}$, add the reaction
@@
We require the following result due to Chen, Doty, Soloveichik [Chen2012DeterministicFunction], guaranteeing that any semilinear function can be built from affine partial functions.
Lemma 5.3 ([Chen2012DeterministicFunction]).
@@
By [Ito1969SemilinearSetsFiniteUnionDisjointLinearSets, Theorem 2], every semilinear set is a finite union of disjoint fundamental linear sets. The author defines a linear set $L=\big{\{}\mathbf{b}+n_{1}\mathbf{u}_{1}+\ldots+n_{p}\mathbf{u}_{p}\mid n_{1},\ldots,n_{p}\in\mathbb{N}\big{\}}$ as fundamental, if $\mathbf{u}_{1},\dots\mathbf{u}_{p}\in\mathbb{N}^{k}$ span a $p$-dimensional vector space in $\mathbb{R}^{k}$, i.e. all vectors are linearly independent in $\mathbb{R}^{k}$.444 This distinction is significant because not all integer-valued linear sets can be represented using solely linearly independent vectors. An illustrative example is $\mathbf{b}=\mathbf{0},\mathbf{u}_{1}=(1,1,1),\mathbf{u}_{2}=(2,0,1),\mathbf{u}_{3}=(0,2,1)$, as discussed in [Chen2012DeterministicFunction]. The vectors $\mathbf{u}_{1},\mathbf{u}_{2},\mathbf{u}_{3}$ are not linearly independent in $\mathbb{R}^{3}$, yet this set cannot be expressed with less than three basis vectors. The proof of Lemma 5.3 in [Chen2012DeterministicFunction] shows that each linear set $L_{i}$ comprising the semilinear graph of $f$ corresponds to one partial affine function $f_{i}$. The fact that [Ito1969SemilinearSetsFiniteUnionDisjointLinearSets, Theorem 2] lets us assume each $L_{i}$ is disjoint from the others immediately implies that each $\mathrm{dom}\ f_{i}$ is disjoint from the others.
The next theorem shows that semilinear functions can be computed by execution bounded CRCs in expected time $\Theta(n\log n)$.
@@
Let $f:\mathbb{N}^{d}\rightarrow\mathbb{N}$ be a semilinear function. Then there is an execution bounded CRC that stably computes $f$ with stabilization time $O(n\log n)$, in expectation and with probability at least $1-n^{-c}$.
@@
We employ the same construction of [Chen2012DeterministicFunction] with minor alterations. A CRC with input species $\Sigma=\{X_{1},\ldots,X_{d}\}$ and output species $\Gamma=\{Y\}$. By Section 5, we decompose our semilinear function into partial affine functions (with linear, disjoint domains), which can be computed in parallel by Lemma 5.1. Further, we decide which function to use by computing the predicate $\phi_{i}=$$\mathrm{x}\in\operatorname{dom}f_{i}$” (Theorem 4.7). We interpret each $\widehat{Y}_{i}^{P}$ and $\widehat{Y}_{i}^{C}$ as an “inactive” version of “active” output species $Y_{i}^{P}$ and $Y_{i}^{C}$. Let $L_{i}^{Y},L_{i}^{N}$ be the yes and no voters respectively voting whether $\mathbf{x}$ lies in the domain of $i$-th partial function. Now, we convert the function result of the applicable partial affine function to the global output by adding the following reactions for each $i\in\{1,\ldots,m\}$.

### v3

@@ 2.4 Time model
The following model of stochastic chemical kinetics is widely used in quantitative biology and other fields dealing with chemical reactions between species present in small counts [10]. It ascribes probabilities to execution sequences, and also defines the time of reactions, allowing us to study the computational complexity of the CRN computation in Sections 4 and 5. If the volume is defined to be the total number of molecules, then the time model is essentially equivalent to the notion of parallel time studied in population protocols [2]. In this paper, the rate constants of all reactions are $1$, and we define the kinetic model with this assumption. A reaction is unimolecular if it has one reactant and bimolecular if it has two reactants. We use no higher-order reactions in this paper.
@@
The kinetic model is based on the physical assumption of well-mixedness that is valid in a dilute solution. Thus, we assume the finite density constraint, which stipulates that a volume required to execute a CRN must be proportional to the maximum molecular count obtained during execution [17]. In other words, the total concentration (molecular count per volume) is bounded. This realistically constrains the speed of the computation achievable by CRNs.
@@ 2.5 Semilinear sets, predicates, functions
The following is a known characterization of the computational power of CRNs [3, 4].
Theorem 2.3 ([3, 4]).
@@ 3 Execution bounded chemical reaction networks
This is a distinct concept from the notion of “bounded” CRNs studied by Rackoff [16] (studied under the equivalent formalism of vector addition systems). That paper defines a CRN to be bounded from a configuration $\mathbf{x}$ if $|\mathsf{reach}(\mathbf{x})|$ is finite (and shows that the decision problem of determining whether this is true is $\mathsf{EXPSPACE}$-complete.) We use the term execution bounded to avoid confusion with this concept.
@@
We first observe that being execution bounded from $\mathbf{x}$ implies a slightly stronger condition: there is a uniform upper bound on the length of all executions from $\mathbf{x}$111 In other words, this rules out the possibility that, although all executions from $\mathbf{x}$ are finite, there are infinitely many of them $\mathcal{E}_{1},\mathcal{E}_{2},\dots$, each finite but longer than the previous.
@@
A CRN is execution bounded from $\mathbf{x}$ if and only if there is a constant $N\in\mathbb{N}$ such that all executions from $\mathbf{x}$ have length at most $N$. Equivalently, there are finitely many executions from $\mathbf{x}$, all of which are finite.
@@
A CRN is execution bounded from $\mathbf{x}$ if and only if $\mathsf{reach}(\mathbf{x})=\{\mathbf{y}\mid\mathbf{x}\Rightarrow\mathbf{y}\}$ is finite and, for all $\mathbf{y}\in\mathsf{reach}(\mathbf{x})$, $\mathbf{y}\not\Rightarrow\mathbf{y}$ except by the zero-length execution.
@@
Every configuration reachable from $\mathbf{x}$ is reached through some execution contained in $T_{\mathcal{C}}(\mathbf{x})$ as a node and there exists only a finite number of them (Section 3). Multiple unique executions can produce the same configuration but one execution cannot produce multiple configurations. Thus, there exists a surjection from the nodes of $T_{\mathcal{C}}(\mathbf{x})$ into $\mathsf{reach}(\mathbf{x})$ and $\mathsf{reach}(\mathbf{x})$ must also be finite. For the second part of the condition, we prove its contrapositive and assume there exists $\mathbf{y}\in\mathsf{reach}(\mathbf{x})$, $\mathbf{y}\Rightarrow_{P}\mathbf{y}\land|P|>0$. Let $P=(\mathbf{p}_{1},\mathbf{p}_{2},\dots,\mathbf{p}_{n})$. It holds that $\mathbf{p}_{1}=\mathbf{p}_{n}$ and $\mathbf{p}_{n-1}\Rightarrow\mathbf{p}_{1}$. We can construct an infinite-length execution $P^{\prime}=(\mathbf{x},\dots,\mathbf{p}_{1},\dots,\mathbf{p}_{n-1},\mathbf{p}_{1},\dots)$, which must also be a valid under the reactions of $\mathcal{C}$, making $\mathcal{C}$ execution unbounded from $\mathbf{x}$.
If $\mathsf{reach}(\mathbf{x})$ is finite and contains no such $\mathbf{y}$, then we can construct a finite, directed, acyclic graph $G_{\mathcal{C}}(\mathbf{x})=(V,E)$ where $V=\mathsf{reach}(\mathbf{x})$, $E=\{(\mathbf{x},\mathbf{y})\mid\mathbf{x}\Rightarrow_{P}\mathbf{y}\land|P|>0\}$. The longest path in the graph has length of at most $|\mathsf{reach}(\mathbf{x})|-1$. A bijection exists between paths in $G_{\mathcal{C}}(\mathbf{x})$ and executions possible in $\mathcal{C}$ starting from $\mathbf{x}$. We set $n=|\mathsf{reach}(\mathbf{x})|$ satisfying that each execution has length of at most $n$, making $\mathcal{C}$ execution bounded.
@@
A execution $\mathcal{E}=(\mathbf{x}_{1},\mathbf{x}_{2},\dots)$ is self-covering if for some $i<j$, $\mathbf{x}_{i}\leqq\mathbf{x}_{j}$. It is strictly self-covering if $\mathbf{x}_{i}\leq\mathbf{x}_{j}$. We also refer to these as (strict) self-covering paths.222 Rackoff [16] uses the term “self-covering” to mean what we call strictly self-covering here, and points out that Karp and Miller [13] showed that $|\mathsf{reach}(\mathbf{x})|$ is infinite if and only if there is a strictly self-covering path from $\mathbf{x}$. The distinction between these concepts is illustrated by the CRN $A\mathop{\rightleftharpoons}\limits B$. From any configuration $\mathbf{x}$, $\mathsf{reach}(\mathbf{x})$ is finite ($|\mathsf{reach}(\mathbf{x})|=\mathbf{x}(A)+\mathbf{x}(B)+1$), and there is no strict self-covering path. However, from (say) $\{A\}$, there is a (nonstrict) self-covering path $\{A\}\Rightarrow\{B\}\Rightarrow\{A\}$, and by repeating, this CRN has an infinite cycling execution within its finite configuration space $\mathsf{reach}(\{A\})=\{\{A\},\{B\}\}$.
@@
For (the contrapositive of) the forward direction, assume there is a self-covering path from $\mathbf{x}$, which reaches to $\mathbf{x}_{i}$ and later to $\mathbf{x}_{j}\geqq\mathbf{x}_{i}$. By additivity the reactions leading from $\mathbf{x}_{i}$ to $\mathbf{x}_{j}$ can be repeated indefinitely (in a cycle if $\mathbf{x}_{i}=\mathbf{x}_{j}$, and increasing some molecular counts unboundedly if $\mathbf{x}_{i}\leq\mathbf{x}_{j}$), so $\mathcal{C}$ is not execution bounded from $\mathbf{x}$.
@@ 4 Execution bounded CRDs stably decide all semilinear sets
In this section, we will show that execution bounded CRDs have the same computational power as unrestricted CRDs. The following is the main result of this section.
@@
Every mod set $M=\big{\{}\left(x_{1},\ldots,x_{d}\right)\mid\sum_{i=1}^{d}w_{i}x_{i}\equiv c\bmod m\big{\}}$ is stably decidable by an execution bounded, single-voting CRD with expected stabilization time $\Theta(n\log n)$.
@@
Every threshold set $T=\big{\{}(x_{1},\ldots,x_{d})\mid\sum_{i=1}^{d}w_{i}x_{i}\geq t\big{\}}$ is stably decidable by an execution bounded, single-voting CRD with expected stabilization time $\Theta(n\log n)$.
@@
If sets $X_{1},X_{2}\subseteq\mathbb{N}^{d}$ are stably decided by some execution bounded, single-voting CRD, then so are $X_{1}\cup X_{2},X_{1}\cap X_{2}$, and $\overline{X_{1}}$ with expected stabilization time $O(n\log n)$.
@@
by two separate CRDs. Subsequently, we add reactions to record the separate votes in one of four new voter species: $V_{NN},V_{NY},V_{YN},V_{YY}$. The first and second CRN determine the first and second subscript respectively. For $b\in\{Y,N\}$ and if $S_{b},T_{b}$ are voters of $\mathcal{C}_{1}$ and $\mathcal{C}_{2}$ respectively, add the reactions:
@@
Above, the ? subscript is shorthand for “any bit”; e.g. if $N_{1}$ is the no voter of the first CRD, we would add two reactions $N_{1}+L_{YN}\rightarrow N_{1}+L_{NN}$ and $N_{1}+L_{YY}\rightarrow N_{1}+L_{NY}$. We let the yes voters be: $\Upsilon=\{V_{NY},V_{YN},V_{YY}\}$ to stably decide $X_{1}\cup X_{2}$ or $\Upsilon=\{V_{YY}\}$ to stably decide $X_{1}\cap X_{2}$.
@@
Since semilinear predicates are exactly Boolean combinations of threshold and mod predicates, Sections 4, 4 and 4.5 imply Theorem 4.1.
@@
Every semilinear set is stably decidable by an execution bounded, all-voting CRD, with expected stabilization time $O(n\log n).$
Proof 4.7.
@@ 5 Execution bounded CRCs stably compute all semilinear functions
Similar to [5, 7], we compute semilinear functions by decomposing them into “affine pieces”, which we will show can be computed by execution bounded CRNs and combined by using semilinear predicates to decide which linear function to apply for a given input.333 While this proof generalizes to multivariate output functions as in [5, 7], to simplify notation we focus on single output functions. Multi-valued functions $f:\mathbb{N}^{d}\to\mathbb{N}^{l}$ can be equivalently thought of as $l$ separate single output functions $f_{i}:\mathbb{N}^{d}\to\mathbb{N}$, which can be computed in parallel by independent CRCs.
@@
We say that a partial function $\hat{f}:\mathbb{N}^{k}\rightarrow\mathbb{N}^{2}$ is a diff-representation of $f$ if $\operatorname{dom}f=\operatorname{dom}\hat{f}$ and, for all $\mathbf{x}\in\operatorname{dom}f$, if $\left(y_{P},y_{C}\right)=\hat{f}(\mathbf{x})$, then $f(\mathbf{x})=y_{P}-y_{C}$, and $y_{P}=O(f(\mathbf{x}))$. In other words, $\hat{f}$ represents $f$ as the difference of its two outputs $y_{P}$ and $y_{C}$, with the larger output $y_{P}$ possibly being larger than the original function’s output, but at most a multiplicative constant larger [7].
@@
Let $f:\mathbb{N}^{k}\rightarrow\mathbb{N}$ be an affine partial function. Then there is a diff-representation $\hat{f}:\mathbb{N}^{k}\longrightarrow\mathbb{N}^{2}$ of $f$ and an execution bounded CRC that monotonically stably computes $\hat{f}$ in expected stabilization time $O(n)$.
@@
For the $c_{i}$ offset, we must reduce the number of $X_{i}$ by $c_{i}$. Since the result will be used in the next reaction, we want to produce a new species $X_{i}^{\prime}$ and require $X_{i}^{\prime}$ to not be consumed during the computation. We achieve this by adding reactions that let $X_{i}$ consume itself $c_{i}$ times (keeping track with a subscript) and converting $X_{i}$ to $X_{i}^{\prime}$ once $c_{i}$ has been reached. For the sake of notation below, assume input species $X_{i}$ is actually named $X_{i,1}$. For each $i\in\{1,\ldots,k\}$ and $m,p\in\left\{1,\ldots,c_{i}\right\}$, if $m+p\leq c_{i}$, add the reaction
@@
We require the following result due to Chen, Doty, Soloveichik [5], guaranteeing that any semilinear function can be built from affine partial functions.
Lemma 5.3 ([5]).
@@
By [12, Theorem 2], every semilinear set is a finite union of disjoint fundamental linear sets. The author defines a linear set $L=\big{\{}\mathbf{b}+n_{1}\mathbf{u}_{1}+\ldots+n_{p}\mathbf{u}_{p}\mid n_{1},\ldots,n_{p}\in\mathbb{N}\big{\}}$ as fundamental, if $\mathbf{u}_{1},\dots\mathbf{u}_{p}\in\mathbb{N}^{k}$ span a $p$-dimensional vector space in $\mathbb{R}^{k}$, i.e. all vectors are linearly independent in $\mathbb{R}^{k}$.444 This distinction is significant because not all integer-valued linear sets can be represented using solely linearly independent vectors. An illustrative example is $\mathbf{b}=\mathbf{0},\mathbf{u}_{1}=(1,1,1),\mathbf{u}_{2}=(2,0,1),\mathbf{u}_{3}=(0,2,1)$, as discussed in [5]. The vectors $\mathbf{u}_{1},\mathbf{u}_{2},\mathbf{u}_{3}$ are not linearly independent in $\mathbb{R}^{3}$, yet this set cannot be expressed with less than three basis vectors. The proof of Lemma 5.3 in [5] shows that each linear set $L_{i}$ comprising the semilinear graph of $f$ corresponds to one partial affine function $f_{i}$. The fact that [12, Theorem 2] lets us assume each $L_{i}$ is disjoint from the others immediately implies that each $\mathrm{dom}\ f_{i}$ is disjoint from the others.
The next theorem shows that semilinear functions can be computed by execution bounded CRCs in expected time $O(n\log n)$.
@@
Let $f:\mathbb{N}^{d}\to\mathbb{N}$ be a semilinear function. Then there is an execution bounded CRC that stably computes $f$ with expected stabilization time $O(n\log n)$.
@@
We employ the same construction of [5] with minor alterations. A CRC with input species $\Sigma=\{X_{1},\ldots,X_{d}\}$ and output species $\Gamma=\{Y\}$. By Section 5, we decompose our semilinear function into partial affine functions (with linear, disjoint domains), which can be computed in parallel by Lemma 5.1. Further, we decide which function to use by computing the predicate $\phi_{i}=$$\mathrm{x}\in\operatorname{dom}f_{i}$” (Theorem 4.1). We interpret each $\widehat{Y}_{i}^{P}$ and $\widehat{Y}_{i}^{C}$ as an “inactive” version of “active” output species $Y_{i}^{P}$ and $Y_{i}^{C}$. Let $L_{i}^{Y},L_{i}^{N}$ be the yes and no voters respectively voting whether $\mathbf{x}$ lies in the domain of $i$-th partial function. Now, we convert the function result of the applicable partial affine function to the global output by adding the following reactions for each $i\in\{1,\ldots,m\}$.
@@
The underlying CRNs computing the predicates and functions have expected stabilization time $O(n\log n).$ Once they have stabilized, the slowest reactions described above are those where a leader ($L_{i}^{Y}$ or $L_{i}^{N}$) must convert all outputs, which also takes expected time $O(n\log n)$ by a coupon collector argument.